MODS1 Commissioning Run 6: UTC 2011 March 23

Observer(s): R. Pogge (OSU)
Telescope Operator: G. Becchetti (LBTO)
AO Support: Remote from Tucson as needed
Telescope Support: J. Little (LBTO)
Instrument Support: T. Shih (LBTO)

Summary

Ice covering the building and roof after last night's storm, otherwise clear and sunny but very cold (-10°C) compared to all of the previous nights, and wind has decreased to about 10m/s (down from gusts up in the 35m/s range last night). As of 11am the show melt equipment is on and J. Little has completed his inspection of the roof. No prognosis yet.

Preparations

Details

All times are UTC unless otherwise noted.

2211 - Still sunny, and -4°C with 70% humidity and winds SW at 7-12m/s. The roof is clear except for the crucial brush seal and the mast and weather station which are still encased in tight ice. Vent doors are still covered with ice and won't likely be able to be opened all night. J. Little judges there to be 50/50 chance of opening tonight. I missed my window to go down with the staff at 3pm if it was a no-go, so I'll stay for the night.

0100 - Beautifully clear as we approach sunset. Still around -4°C and slowly falling, humiditys steady at around 70%, wind 10-15m/s out the SW. The roof brush seal is clear of ice, but still considerable rime ice on the weather station mast and lightning ball. The rest of the ice up there is frozen to the roof. The problem is the masts, they constitute an ice risk to the telescope. The ice on the masts is concentrated on the ball and parts pf the weather station mast with no heat tape. John whacked at it with a shovel and it didn't budge.

John's assessment is that we should not open because of the ice risk due to the remaining ice on the masts. John will consult with folks downtown, but this is the right thing to do.

0208 - Sunset, winds SW at 14m/s, -5°C and falling, humidity 65% and slowly falling. Thin bands of high cloud visible way off to the west but clear and velvet dark overhead. Nice view of Jupiter and Mercury in the western sky.

0500 - Temp up to -2.5°C, humidity dropped to 30% in the past hour. Wind is SW 10-15m/s and steady the past hour. It has to remain below 8m/s to open just the left door according to John Hill. We're not there yet.

0915 - We're going to call the night a loss at this point since we cannot safely open.

-- RichardPogge - 22 Mar 2011
Topic revision: r6 - 23 Mar 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback